OpenStudy (anonymous):

Find the sum of all possible values of θ (in degrees) restricted to the domain 0≤θ≤360 ∘ , satisfying: sin^4θ+sin^2θcos^2θ+cos^4θ=3/4 .

OpenStudy (anonymous):

so it should be sin^4 of theta + sin^2 of theta times cosine^2 of theta + cosine^4 of theta = 3/4

OpenStudy (ankit042):

the equation can be written as (sin^2 thea +cos^2 theta)^2-(sin^2theta)(cos^2thea)=3/4

OpenStudy (ankit042):

Now use sin^2 +cos^2 =1

OpenStudy (anonymous):

I don know trig well, i just wanted to know the answer please :)

OpenStudy (ankit042):

I can't give the answer directly you have to try

OpenStudy (jdoe0001):

so just the answer because you dunno trig and somehow that'll help know trig better?

OpenStudy (anonymous):

well with a solution, i haven't figured it out

OpenStudy (anonymous):

I really need help please

OpenStudy (anonymous):

YEAHHHH ITS 720! flutter YOU ALL! BURN IN flutterING HELL! MEXICAN, JEWS, CHINESE KUNICHIWA'S CAN ALL SUCK A sleek-feathered oneS wingspan featherES!